# DATABench: Evaluating Dataset Auditing in Deep Learning from an Adversarial Perspective

### v2

@@ 3 Attack Methodology
As introduced in Section 2.4, robustness and distinctiveness are two fundamental properties of a practical dataset auditing method. However, there is still a significant gap between the constrained settings of existing works and practice for evaluating these two properties. Existing research [51, 20, 33] primarily considers common yet simple operations in DL and lacks dedicated attack evaluation. To bridge this gap, we propose a unified and comprehensive formulation and methodology of attacks against dataset auditing. We also present systematic attack strategies and implementations of these attacks.
@@ 3.2 Evasion Attacks
Attack Strategy. According to Section 2.3, existing dataset auditing methods can be classified into internal feature (IF) based methods and external feature (EF) based methods. This taxonomy can facilitate the design of attacks, as follows.
@@
Attack Implementations. Based on the aforementioned attack strategies, we can design various attacks in different attack phases. The implementations of evasion attacks in DATABench are as follows.
@@
  • Decoupling: To reduce the association between the trained model and the dataset, the adversary can adopt operations to make the samples in the dataset less similar to the original ones. As such, data augmentation and data synthesis are two approaches to achieve such a goal. Data augmentation applies transformations such as cropping and flipping to increase the diversity of training data [78]. Data synthesis is an advanced technique that utilizes the capability of generative models [49, 18]. It first trains a generative model ($e.g.$, a diffusion model [37]) on the dataset and then uses the model to generate a synthetic dataset. The model is trained on the synthetic dataset instead of the original one.
  • Removal: Since the external features usually tend to be ‘noise’ in the dataset, denoising techniques can be adopted to remove them. We consider two classes of denoising methods, traditional denoising filters and DL-based denoising methods. The former reduces noise by smoothing pixel values [26] and the latter exploits DL models ($e.g.$, autoencoder [50]) to remove the noise while preserving the important features for model training.
@@
  • Decoupling: The adversary can reduce the model’s overfitting during training to weaken its association with the training dataset. Regularization is a common approach in practice to achieve such a goal. Additionally, during training, the adversary can apply Differential Privacy (DP) mechanisms [24] ($e.g.$, DP-SGD [1]) to make any single data point indistinguishable.
  • Removal: To eliminate potential abnormal behaviors during training, we consider adversarial training and training-phase backdoor defense as two representative removal attacks. Adversarial training [80] improves the prediction robustness of the model by including adversarial examples (AE) during the training process. AEs are inputs intentionally perturbed to lead to misclassification. Training-phase backdoor defense [28, 41] aims to mitigate hidden external features in the training data by trying not to learn from them. Both ways may remove the designated behaviors and evade dataset auditing.
@@
  • Decoupling: In this phase, the adversary can also add noise during the model inference. As such, we implement noisy output [20] and noisy feature [21] that perturb the outputs and intermediate results of the model, respectively, during inference to evade dataset auditing.
  • Removal: Since the model parameters are fixed, the adversary cannot remove the external features from the model. However, it can still hide the external behaviors by changing inference mechanisms. Randomized smoothing (RS) and transformation-based backdoor defense are two typical methods. Specifically, RS [16] adds noise to the inputs and calculates the aggregated results of the outputs on noisy inputs. Transformation-based backdoor defense [53] ($e.g.$, reprogramming [13]) neutralizes the designated behaviors by applying input transformations.
  • Detection: This type of attack aims to detect the trigger samples used for dataset auditing. Since the trigger samples tend to be different from normal samples, we implement traditional outlier detection [6] based on the features of inputs and advanced backdoor sample detection [38, 32]. If these methods identify the input samples as an outlier or backdoor sample, the adversary can refuse to reply to evade dataset auditing.
@@ 3.3 Forgery Attacks
Attack Strategy. The goal of the forgery attack is to craft samples to falsely claim that an independent model $\mathcal{M}_{u}$ is trained on these samples. Assuming that the adversary has an auxiliary dataset $\mathcal{D}_{a}$, they can optimize a set of perturbations $\mathcal{R}=\{\bm{r}_{i}\}_{i}^{|\mathcal{D}_{a}|}$ to make the dataset auditing method output $1$. Ideally, this can be formalized as an optimization problem:
@@
where $(\bm{x}_{i},\bm{y}_{i}),\bm{r}_{i}$ are the $i$-th element in $\mathcal{D}_{a}$ and $\mathcal{R}$. Inspired by [57], the adversary can adopt adversarial attack techniques [29] to achieve the forgery attack.
@@
$\mathcal{T}(\mathcal{M},\bm{x},\bm{y}_{t})=\arg\min_{\bm{r}}\mathcal{L}(\mathcal{M}(\bm{x}+\bm{r}),\bm{y}_{t}),\ \text{s.t.}\ \|\bm{r}\|\leq\epsilon,$ (7)
@@
Data Synthesis Evasion Decoupling Pre-processing
Gaussian Filter, Median Filter,
Wavelet Filter
Evasion Removal Pre-processing
Autoencoder-based Denoising Evasion Removal Pre-processing
Regularization, DP-SGD [1] Evasion Decouping Training
Adversarial Training [80] Evasion Removal Training
Adaptively Splitting Poisoned
Dataset (ASD) [28]
Evasion Removal Training
Noisy Output, Noisy Feature Evasion Decoupling Post-training
Randomized Smoothing (RS) [16] Evasion Removal Post-training
Reprogramming Evasion Removal Post-training
SCALE-UP [32] Evasion Detection Post-training
KNN-based Outlier Detection,
SVM-based Outlier Detection
Evasion Detection Post-training
FGSM [29], PGD [59], UAP [64]
TIFGSM [19], VNIFGSM [82]
Forgery / /
@@
4 DATABench: Benchmarking Attacks against Dataset Auditing
4.1 Overview of DATABench
In this section, we provide an overview of our dataset auditing attack benchmark, DATABench. This work primarily focuses on dataset auditing for image classification models, as most existing studies concentrate on this domain and cannot be readily extended to other modalities or tasks. A discussion on dataset auditing for generative models is deferred to Section 5.2. In particular, our DATABench is designed to encompass both basic and advanced attack strategies. Table 1 shows the implemented attacks in DATABench. We implement a total of 17 evasion attacks, including 6 attacks in the pre-processing phase, 4 attacks in the training phase, and 7 attacks in the post-training phase. We exclude naive strategies such as pruning and quantization, as prior studies have consistently shown their ineffectiveness [51, 20, 42, 54]. Besides, we implement 5 different forgery attacks. Detailed descriptions of these attacks are in Appendix B.
In DATABench, we implement 9 representative dataset auditing methods, including 4 IF-based methods and 5 EF-based methods. For IF-based methods, besides Dataset Inference (DI) [61] and Data-use Auditing (DUA) [42], which are designed for dataset auditing, we also adopt two representative MIAs as auditing methods. We select a classic MIA proposed by Shokri et al. [77] (denoted as MIA) and the state-of-the-art (SOTA) MIA named Rapid [36]. For EF-based methods, we evaluate DVBW [54], UBW (including UBW-P and UBW-C) [51], ZeroMark [31], and Domain Watermark (DW) [33]. The detailed settings can be found in Appendix A.
@@ 4.2 Experimental Settings
Models and Datasets. In the main experiments, we adopt the widely used CIFAR-10 [47] dataset and the ResNet-18 [34] architecture. This setting, although simple from an auditing perspective, presents substantial challenges for adversaries, making it well-suited for identifying vulnerabilities of existing methods. To further demonstrate the generalizability of our findings, we also conduct experiments on other models and datasets, including Vision Transformer [62] and ImageNet [17], in Appendix G. We further split the training data: 70% is the public release dataset (protected), 15% is the defender’s unpublished audit auxiliary dataset, and 15% is the adversary’s unpublished attack auxiliary dataset. Both auxiliary sets are assumed to be trusted and clean (unwatermarked). We train the model from scratch on the victim dataset. The experiments in the fine-tuning scenario are in Appendix H.
Attack Settings. In our experiments, we evaluate existing dataset auditing methods against all attacks included in DATABench. Since data augmentation and regularization substantially affect model utility, we adopt them as default settings. For adversarial training attacks, we consider two variants: AdvTraining-FGSM, which leverages FGSM [29] to generate adversarial examples, and AdvTraining-Hybrid, which combines three attack methods [29, 59, 8]. For DP-SGD, we examine two privacy budgets ($\varepsilon=32$ and $64$) to simulate different levels of privacy protection. Detailed descriptions and configurations are provided in Appendix B. To ensure both simplicity and fairness, we adopt the most classical or their default configurations for each attack. We observe that this might introduce model utility degradation in certain scenarios, which we further discuss in Appendix E.
@@
Metrics. We use the accuracy on the testing dataset (Test Acc.) to evaluate the utility of the attacked models. Different metrics are applied to evaluate the effectiveness of different dataset auditing methods, as follows.

### v3

@@ 4 Attack Methodology
As introduced in Section 3.3, robustness and distinctiveness are two fundamental properties of a practical dataset auditing method. However, there is still a significant gap between the constrained settings of existing works and practice for evaluating these two properties. Existing dataset auditing research [li2022untargeted, du2025sok, guo2023domain] primarily considers common yet simple operations in DL and lacks dedicated attack evaluation. To bridge this gap, we propose a unified and comprehensive formulation and methodology of attacks against dataset auditing. We also present systematic attack strategies and implementations of these attacks. All these elements lead to the development of our Dataset Auditing Attack Benchmark, DATABench.
@@ 4.2 Evasion Attacks
Attack Strategy. Based on Section 3.2, existing dataset auditing methods can be classified into internal feature (IF) based methods and external feature (EF) based methods. This taxonomy can facilitate the design of attacks, as follows.
@@
Attack Implementations. Based on the attack strategies, we can design various attacks in different attack phases. The implementations of evasion attacks in DATABench are as follows.
@@
  • Decoupling: To reduce the association between the trained model and the dataset, the adversary can adopt operations to make the samples in the dataset less similar to the original ones. As such, data augmentation and data synthesis are two approaches to achieve such a goal. Data augmentation applies transformations such as cropping and flipping to increase the diversity of training data [shorten2019survey]. Data synthesis is an advanced technique that utilizes the capability of generative models [li2025easy, dockhorn2023differentially]. It first trains a generative model ($e.g.$, a diffusion model [ho2020denoising]) on the dataset and then uses the model to generate a synthetic dataset. The model is subsequently trained on the synthetic dataset instead of the original one.
  • Removal: Since the external features usually tend to be ‘noise’ in the dataset, denoising techniques can be adopted to remove them. We consider two classes of denoising methods, traditional denoising filters and DL-based denoising methods. The former reduces noise by smoothing pixel values [fan2019brief] and the latter exploits DL models ($e.g.$, autoencoder [li2023comprehensive]) to remove the noise while preserving the important features for model training.
@@
  • Decoupling: The adversary can reduce the model’s overfitting during training to weaken its association with the training dataset. For instance, regularization is a common approach in practice to achieve such a goal. Additionally, during training, the adversary can apply Differential Privacy (DP) mechanisms [dwork2006calibrating] ($e.g.$, DP-SGD [abadi2016deep]) to make any single data point indistinguishable.
  • Removal: To eliminate potential abnormal behaviors during training, we consider adversarial training and training-phase backdoor defense as two representative removal attacks. Adversarial training [tramer2018ensemble] improves the prediction robustness of the model by including adversarial examples (AE) during the training process. AEs are inputs intentionally perturbed to lead to misclassification; Training-phase backdoor defense [gao2023backdoor, huang2022backdoor] aims to mitigate hidden external features in the training data by trying not to learn from them. Both ways may remove the designated behaviors and evade dataset auditing.
@@
  • Decoupling: In this phase, the adversary can also apply DP mechanisms to add noise during the model inference. As such, we implement noisy output [du2025sok] and noisy feature [du2023dp] that perturb the outputs and intermediate results of the model, respectively, during inference to evade dataset auditing.
  • Removal: Since the model parameters are fixed, the adversary cannot remove the external features from the model. However, it can still hide the external behaviors by changing inference mechanisms. Randomized smoothing (RS) and transformation-based backdoor defense are two typical methods. Specifically, RS [cohen2019certified] adds noise to the inputs and calculates the aggregated results of the outputs on noisy inputs; Transformation-based backdoor defense [li2021backdoor] ($e.g.$, reprogramming [chen2025refine]) neutralizes the designated behaviors by applying input transformations.
  • Detection: This type of attack aims to detect the trigger samples used for dataset auditing. Since the trigger samples tend to be different from normal samples, we implement traditional outlier detection [boser1992training] based on the features of inputs and advanced backdoor sample detection [hou2024ibd, guo2023scale]. If these methods identify the input samples as an outlier or backdoor sample, the adversary can refuse to reply to evade dataset auditing.
@@ 4.3 Forgery Attacks
Attack Strategy. The goal of the forgery attack is to craft some samples to falsely claim that an independent model $\mathcal{M}_{u}$ is trained on these samples. Assuming that the adversary has an auxiliary dataset $\mathcal{D}_{a}$, he can optimize a set of perturbations $\mathcal{R}=\{\bm{r}_{i}\}_{i}^{|\mathcal{D}_{a}|}$ to make the dataset auditing method output $1$. Ideally, this can be formalized as an optimization problem:
@@
where $(\bm{x}_{i},\bm{y}_{i}),\bm{r}_{i}$ are the $i$-th element in $\mathcal{D}_{a}$ and $\mathcal{R}$. Inspired by [liu2024false, shao2025fit], the adversary could adopt adversarial attack techniques [goodfellow2014explaining, ren2020adversarial] to achieve the forgery attack.
@@
$\mathcal{T}(\mathcal{M},\bm{x},\bm{y}_{t})=$ $\arg\min_{\bm{r}}\mathcal{L}(\mathcal{M}(\bm{x}+\bm{r}),\bm{y}_{t}),$ (7)
s.t. $\|\bm{r}\|\leq\epsilon,$
@@
Data Synthesis [ho2020denoising] Evasion Decoupling Pre-processing
Gaussian Filter, Median Filter, Wavelet Filter [fan2019brief] Evasion Removal Pre-processing
Autoencoder-based Denoising [li2023comprehensive] Evasion Removal Pre-processing
Regularization, DP-SGD [abadi2016deep] Evasion Decouping Training
Adversarial Training [tramer2018ensemble] Evasion Removal Training
Adaptively Splitting Poisoned Dataset (ASD) [gao2023backdoor] Evasion Removal Training
Noisy Output, Noisy Feature [du2025sok] Evasion Decoupling Post-training
Randomized Smoothing (RS) [cohen2019certified] Evasion Removal Post-training
Reprogramming [chen2025refine] Evasion Removal Post-training
SCALE-UP [guo2023scale] Evasion Detection Post-training
KNN-based Outlier Detection, SVM-based Outlier Detection Evasion Detection Post-training
FGSM [goodfellow2014explaining], PGD [madry2018towards], UAP [moosavi2017universal], TIFGSM [dong2019evading], VNIFGSM [wang2021enhancing] Forgery / /
5 Benchmarking Attacks against Dataset Auditing
5.1 Overview of DATABench
In this section, we present the overview of our dataset auditing attack benchmark DATABench. Notably, DATABench aims to cover both simple and advanced attack methods. Table 1 shows the implemented attacks in DATABench. We implement a total of 17 evasion attacks, including 6 attacks in the pre-processing phase, 4 attacks in the training phase, and 7 attacks in the post-training phase. Naive attacks such as pruning and quantization are not included since a wide range of papers have demonstrated they are ineffective [li2022untargeted, du2025sok, huang2024general, li2023black]. Besides, we implement 5 different forgery attacks. Detailed descriptions of these attacks are in Appendix B.
In DATABench, we implement 9 representative dataset auditing methods, including 4 IF-based methods and 5 EF-based methods. For IF-based methods, besides Dataset Inference (DI) [maini2020dataset] and Data-use Auditing (DUA) [huang2024general], which are designed for dataset auditing, we also adopt two representative MIAs as auditing methods. We select a classic MIA proposed by Shokri et al. [shokri2017membership] (denoted as MIA) and the state-of-the-art (SOTA) MIA named Rapid [he2024difficulty]. For EF-based methods, we evaluate DVBW [li2023black], UBW (including UBW-P and UBW-C) [li2022untargeted], ZeroMark [guo2024zeromark], and Domain Watermark (DW) [guo2023domain]. The detailed settings of these dataset auditing methods can be found in Appendix A.
@@ 5.2 Experimental Settings
Models and Datasets. In the main experiments, we adopt the widely-used CIFAR-10 [krizhevsky2009learning] and ResNet-18 [he2016deep]. The reason for using this simple model and dataset is that almost all the existing dataset copyright auditing methods perform the best in a simple dataset [du2025sok]. We also include experiments on other models and datasets such as Vision Transformer [mehta2021mobilevit] and ImageNet [deng2009imagenet] in Section 5.3. We further split the training data: 70% is the public release dataset (protected), 15% is the defender’s unpublished audit auxiliary dataset, and 15% is the adversary’s unpublished attack auxiliary dataset. Both auxiliary sets are assumed to be trusted and clean (unwatermarked). We train the model both from scratch on the victim dataset and by fine-tuning (Section 5.4).
Attack Settings. In our experiments, we evaluate existing dataset auditing methods on all the attacks in DATABench. Given the significant impact of data augmentation and regularization on model utility, we adopt them as default settings. Besides, we consider two variants for adversarial training attacks. ‘AdvTraining-FGSM’ simply uses FGSM [goodfellow2014explaining] to generate the adversarial examples for training, and ‘AdvTraining-Hybrid’ composes three different adversarial attacks [goodfellow2014explaining, madry2018towards, carlini2017towards]. We also test two different privacy budgets ($\varepsilon=32$ or $64$) in DP-SGD to simulate different levels of privacy protection. Detailed descriptions and settings can be found in Appendix B.
@@
Metrics. We use the accuracy on the testing dataset (Test Acc.) to evaluate the utility of the attacked models. We mark cases where the test accuracy drops significantly ($10\%$) as failed attacks. Different metrics are applied to evaluate the effectiveness of different dataset auditing methods, as follows.